OpenStudy (anonymous):

integral from 0 to 1 for lnx/x dx

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

Since \[\frac{lnx}{x}\] doesn't seem to have an elementary anti-derivative, we have to mess around with it. The first thing I always see is if there is a u-substitution that turns it into something I know. Any ideas what you could substitute?

OpenStudy (anonymous):

yah so we would let u=lnx du=1/x dx

OpenStudy (anonymous):

then we have integral of 0 to 1 of u

OpenStudy (anonymous):

Then what would that make your integral look like?

OpenStudy (anonymous):

then we have u^2/2 but im lost from here i know it is divergent but i dont know how to show that

OpenStudy (anonymous):

We have \[\frac{u^2}{2} + c\] But it isn't from 0 to 1. Because\[u = lnx, x=0, u = -\infty, x=1, u = 0\] If we use these limits, what does that say about the divergence?

OpenStudy (anonymous):

well so, im lost sorry so what im doing is [ln^2(1)/2 - ln^2(0)/2] and then it is [0-und.] so it makes it undefined?

OpenStudy (anonymous):

this is plugging u=lnx back into antideriv

OpenStudy (anonymous):

If you have any discontinuities in the function, in this case \(\ln(x)/x\) is not continuous at \(x=0\), then you have to use a limit to evaluate it.

OpenStudy (anonymous):

ok so i would use the lim as x aproachs 0 ?

OpenStudy (anonymous):

A one sided limit, depending on which integration limit we are talking about.

OpenStudy (anonymous):

Since 0 is a lower limit, we would do \(x\to 0^{-}\)

OpenStudy (anonymous):

This is the part of calculus that I haven't touched in eight years, a bit rusty on it.

OpenStudy (anonymous):

ok im going to do so its ok as atleast you tried helping thats really nice!

OpenStudy (anonymous):

\[ \lim_{x\to 0^-}\frac{\ln^2(1) -\ln^2(x)}{2} \]

OpenStudy (anonymous):

ok so would i do lim as x aproaches 0^- integral of x to 1 of lnx/x

OpenStudy (anonymous):

It simplifies to \[ \dfrac{0-\lim_{x\to 0^-}\ln^2(x)}{2} \]

OpenStudy (anonymous):

this graph doesnt exists from 0^- direction

OpenStudy (anonymous):

hmm, actually maybe I should have said \(0^+\)

OpenStudy (anonymous):

ok that i think is better cause the graph doesnt exists its domain is (0,infin)

OpenStudy (anonymous):

In either case, this integral results in \(-\infty\)

OpenStudy (anonymous):

isnt it positive infinity?

OpenStudy (anonymous):

\[ \frac{0-\infty}{2} \to -\infty \]

OpenStudy (anonymous):

ok thnks